# Disentangle Identity, Cooperate Emotion: Correlation-Aware Emotional Talking Portrait Generation

### v1

@@ Abstract.
Recent advances in Talking Head Generation (THG) have achieved impressive lip synchronization and visual quality through diffusion models; yet existing methods struggle to generate emotionally expressive portraits while preserving speaker identity. We identify three critical limitations in current emotional talking head generation: insufficient utilization of audio’s inherent emotional cues, identity leakage in emotion representations, and isolated learning of emotion correlations. To address these challenges, we propose a novel framework dubbed as DICE-Talk, following the idea of disentangling identity with emotion, and then cooperating emotions with similar characteristics. First, we develop a disentangled emotion embedder that jointly models audio-visual emotional cues through cross-modal attention, representing emotions as identity-agnostic Gaussian distributions. Second, we introduce a correlation-enhanced emotion conditioning module with learnable emotion banks that explicitly capture inter-emotion relationships through vector quantization and attention-based feature aggregation. Third, we design an emotion discrimination objective that enforces affective consistency during the diffusion process through latent-space classification. Extensive experiments on MEAD and HDTF datasets demonstrate our method’s superiority, outperforming state-of-the-art approaches in emotion accuracy while maintaining competitive lip-sync performance. Qualitative results and user studies further confirm our method’s ability to generate identity-preserving portraits with rich, correlated emotional expressions that naturally adapt to unseen identities. Code is available at GitHub.
Talking head generation, emotion control, diffusion model
copyright: none
Figure 1. Our DICE-Talk, as the first emotional THG method based on diffusion models, can generate various emotions while well preserving the identity characteristic, thus largely benefiting the real-life application of digital human.
@@ 1. Introduction
Recent advancements in generative models have shed light on generating high-quality and realistic videos under various controlling conditions such as texts (Rombach et al., 2022), images (Blattmann et al., 2023a), videos (Wu et al., 2023), etc. Among all the different subtasks of video generation, Talking Head Generation (THG), as a human-centric task which aims to generate videos of talking heads guided by conditions such as speech and images, has emerged as a significant problem due to its wide application in scenarios such as digital humans, film production, virtual reality. Despite importance, the extreme challenges result from its low tolerance to artifacts in general and its demand of high fidelity in lip shapes, facial expressions, and head motions.
Following the commonly used generative models, GAN-based THG methods (Prajwal et al., 2020; Zhang et al., 2023) have achieved remarkable results in generating high-resolution videos through adversarial training between generators and discriminators, particularly in terms of visual quality and lip-sync accuracy. Diffusion model-based THG methods (Shen et al., 2023; Tian et al., 2024), on the other hand, excel in generating high-quality and high-resolution images and videos, and it outperforms GANs in terms of the stability and consistency of the generated content, thus becoming the mainstream methods for THG. These methods largely facilitate THG by strengthening the explicit controlling conditions such as facial keypoints and head motion sequences. Yet, a critical gap persists: none of these methods holistically model emotion as a controllable, multimodal signal. Ignoring such a condition forces models to generate ”emotionally flat” outputs, limiting their applicability in fields like mental health support, where emotional resonance is critical.
@@
Overall, our contributions are summarized as follows:
  • We propose DICE-Talk, the first audio-driven diffusion-based THG framework that disentangles speaker identity from emotional expressions. This enables realistic talking head generation with rich emotional expressions while preserving identity fidelity.
  • We introduce a cross-modal disentangled emotion embedder, representing emotions as identity-agnostic Gaussian distributions to address identity leakage and leverage audio’s inherent prosodic cues.
  • A correlation-enhanced emotion conditioning mechanism is proposed, which leverages a learnable emotion bank to capture inter-emotion relationships.
@@ 2. Related Work
Diffusion Model-Based Talking Head Generation. Recently, there has been a surge of research  (Shen et al., 2023; Xie et al., 2024; Tian et al., 2024; Wang et al., 2024; Wei et al., 2024; Xu et al., 2024; Yang et al., 2024; Cui et al., 2024; Ji et al., 2024; Jiang et al., 2024) utilizing diffusion models to achieve high-quality portrait videos. Among these, X-Portrait (Xie et al., 2024) and MegActor (Yang et al., 2024) rely on the pose and expression from the source video to generate the target video, which limits their ability to produce videos based solely on audio. DiffTalk (Shen et al., 2023) was the first to modify lip movements using audio and diffusion models, but it does not extend to driving other head parts. EMO (Tian et al., 2024) was the first to leverage LDM (Rombach et al., 2022) and audio features to achieve overall motion in portraits. V-Express (Wang et al., 2024) controls the overall motion amplitude by adjusting audio attention weights, while Hallo (Xu et al., 2024) designed a hierarchical module to regulate the motion amplitude of different regions. Sonic (Ji et al., 2024) focus on global audio perception to obtain better talking head results. In summary, current audio-driven diffusion model approaches have not taken into account that each portrait should exhibit a corresponding emotion while speaking, which is essential for generating higher-quality portrait videos.
Emotional Talking Head Generation. Previous research has explored several GAN-based methods  (Wang et al., 2020; Wu et al., 2021; Ji et al., 2021, 2022; Liang et al., 2022; Sinha et al., 2022; Ma et al., 2023; Gan et al., 2023) for extracting style information to apply in talking head generation. MEAD (Wang et al., 2020) and Emotion (Sinha et al., 2022) directly inject style labels into the network to drive the corresponding emotions. GC-AVT (Liang et al., 2022) and EAMM (Ji et al., 2022) map the facial expressions of each frame in the source video to each frame in the target video. LSF (Wu et al., 2021) and StyleTalk (Ma et al., 2023) employ 3D Morphable Models (3DMM) to extract facial information and construct style codes that drive the desired styles. EAT (Gan et al., 2023) achieves emotion control through parameter-efficient adaptations. Building on these approaches, our framework is first to disentangle identity and cooperate emotion to generate better emotion features into diffusion model and produce high-quality emotional talking portrait videos.
@@ 3. Method
Problem Formulation.  The goal of THG is to generate a talking head video under the control of a reference portrait image, audio, and emotion prior. Among these conditions, the reference portrait image provides the background and facial identity, the audio guides the overall head motion and lip movements. In addition, we propose a disentangled emotion embedder to extract the emotion prior from the visual and audio content of an emotion reference video, which is used to determine facial emotion and speaking habit.
@@
Overview.  As depicted in Figure 2, our DICE-Talk consists of three important designs: (1) A disentangled emotion embedder that extracts emotions from video and audio inputs, representing them as a Gaussian distribution. (2) A correlation-enhanced emotion conditioning module that stores emotion features in a bank and injects them into a diffusion model. (3) An emotion discrimination objective that improves emotion expression by classifying emotions during denoising. The diffusion process combines a reference image, audio features, emotion data, and the emotion bank to guide facial animation. Audio features come from Whisper-Tiny (Radford et al., 2023), and we use an attention layer for processing following Ji et. al. (Ji et al., 2024). Finally, after denoising, a VAE decoder generates expressive portrait frames with the target emotion.
@@ 3.1. Disentangled Emotion Embedder
where $\tau$ denotes a temperature parameter, $\mathcal{S}^{n}$ denotes all negative samples for $s$, and $\zeta(s_{i},s_{j})$ is the cosine similarity between sample pairs $(s_{i},s_{j})$. We additionally add a fixed constant to stabilize the numerical range of the similarity and make the training process more stable.
@@ 3.2. Correlation-Enhanced Emotion Conditioning
Emotion Projection.  To utilize the emotion feature $E_{s}$ obtained in Sec. 3.2 and guide the denoising process, then $E_{s}$ is injected into the diffusion UNet through an additional emotion attention layer, where it interacts with other features via a cross-attention mechanism to supplement additional facial details such as expressions and speaking habits. The information of the emotion feature $E_{s}$ can be injected into the spatial and temporal cross-attention layers to provide spatial knowledge as follows:
@@ 4.2. Quantitative Comparison
We compare our method with several previous works, including SadTalker (Zhang et al., 2023), AniPortrait (Wei et al., 2024), V-Express (Wang et al., 2024), Hallo (Xu et al., 2024), Hallo2 (Cui et al., 2024), EchoMinic (Chen et al., 2024), Sonic (Ji et al., 2024), EAMM (Ji et al., 2022), StyleTalk (Ma et al., 2023), and EAT (Gan et al., 2023). Among these, EAMM, StyleTalk, and EAT are also focused on emotion modeling.
@@
As shown in Table 1, our method ranks among the top in metrics on both the MEAD and out-of-domain datasets. Our method achieved the best emotion scores, reflecting that our approach generates vivid and accurate facial expressions through emotion priors. In the quality assessment of videos, single-frame images, and lip-sync accuracy, our method is close to the current best methods, as evidenced by the FVD, FID, and SyncNet scores. The results in Table 2 further indicate that our method maintains high visual quality even on datasets like HDTF, which do not have obvious emotions. These results demonstrate that our method can ensure video generation quality while achieving diversity and accuracy in emotional expression.
@@ 4.4. User Study
We conducted a subjective evaluation of the results on open datasets, assessing the comparative methods along four dimensions: emotional expression, lip-sync accuracy, video quality, and video smoothness. 20 participants rated the results of five comparative methods on a scale of 1 to 5 in a total of 40 video sets. As shown in Table 3, our DICE-Talk significantly outperformed others in the dimensions of emotional expression and video smoothness, with a notable 66.5% improvement in emotional expression. Although StyleTalk and EAT achieved good results in objective metrics, they compromised the original identity and reduced visual quality while generating emotions, leading to lower subjective emotion scores. Figure 4(a) illustrates two clear examples. In terms of lip-sync accuracy and video quality, our method is very close to the state-of-the-art Sonic and far surpasses other methods that model emotions. Several video results from the user study are included in the supplementary materials.
4.5. Ablation Study

### v2

@@ Abstract.
Recent advances in Talking Head Generation (THG) have achieved impressive lip synchronization and visual quality through diffusion models; yet existing methods struggle to generate emotionally expressive portraits while preserving speaker identity. We identify three critical limitations in current emotional talking head generation: insufficient utilization of audio’s inherent emotional cues, identity leakage in emotion representations, and isolated learning of emotion correlations. To address these challenges, we propose a novel framework dubbed as DICE-Talk, following the idea of disentangling identity with emotion, and then cooperating emotions with similar characteristics. First, we develop a disentangled emotion embedder that jointly models audio-visual emotional cues through cross-modal attention, representing emotions as identity-agnostic Gaussian distributions. Second, we introduce a correlation-enhanced emotion conditioning module with learnable emotion banks that explicitly capture inter-emotion relationships through vector quantization and attention-based feature aggregation. Third, we design an emotion discrimination objective that enforces affective consistency during the diffusion process through latent-space classification. Extensive experiments on MEAD and HDTF datasets demonstrate our method’s superiority, outperforming state-of-the-art approaches in emotion accuracy while maintaining competitive lip-sync performance. Qualitative results and user studies further confirm our method’s ability to generate identity-preserving portraits with rich, correlated emotional expressions that naturally adapt to unseen identities.
Talking head generation; emotion control; diffusion model
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755286isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision tasks
Figure 1. Our DICE-Talk, as an emotional THG method based on diffusion models, can generate various emotions while well preserving the identity characteristic, thus largely benefiting the real-life application of digital human.
@@ 1. Introduction
Recent advancements in generative models have shed light on generating high-quality and realistic videos under various controlling conditions such as texts (Rombach et al., 2022), images (Blattmann et al., 2023a; Huang et al., 2024; Pan et al., 2025), videos (Wu et al., 2023), etc. Among all the different subtasks of video generation, Talking Head Generation (THG), as a human-centric task which aims to generate videos of talking heads guided by conditions such as speech and images, has emerged as a significant problem due to its wide application in scenarios such as digital humans, film production, virtual reality. Despite importance, the extreme challenges result from its low tolerance to artifacts in general and its demand of high fidelity in lip shapes, facial expressions, and head motions.
Following the commonly used generative models, GAN-based THG methods (Prajwal et al., 2020; Zhang et al., 2023) have achieved remarkable results in generating high-resolution videos through adversarial training between generators and discriminators, particularly in terms of visual quality and lip-sync accuracy. Diffusion model-based THG methods (Shen et al., 2023; Tian et al., 2024), on the other hand, excel in generating high-quality and high-resolution images and videos, and outperform the GANs in terms of stability and consistency, thus becoming the mainstream methods for THG. These methods largely facilitate THG by strengthening the explicit controlling conditions such as facial keypoints and head motion sequences. Yet, a critical gap persists: none of these methods holistically model emotion as a controllable, multimodal signal. Ignoring such a condition forces models to generate ”emotionally flat” outputs, limiting their applicability in fields like mental health support, where emotional resonance is critical.
@@
Overall, our contributions are summarized as follows: (1) We propose DICE-Talk, an audio-driven diffusion-based THG framework that disentangles speaker identity from emotional expressions. This enables realistic talking head generation with rich emotions while preserving identity fidelity. (2) We introduce a cross-modal disentangled emotion embedder, representing emotions as identity-agnostic Gaussian distributions to address identity leakage and leverage audio’s inherent prosodic cues. (3) A correlation-enhanced emotion conditioning mechanism is proposed, which leverages a learnable emotion bank to capture inter-emotion relationships.
@@ 2. Related Work
Diffusion Model-Based Talking Head Generation. Recently, there has been a surge of research  (Shen et al., 2023; Xie et al., 2024; Wei et al., 2024; Yang et al., 2024; Cui et al., 2024; Ji et al., 2024; Jiang et al., 2024) utilizing diffusion models to achieve high-quality portrait videos. Among these, X-Portrait (Xie et al., 2024) and MegActor (Yang et al., 2024) rely on the pose and expression from the source video to generate the target video, which limits their ability to produce videos based solely on audio. DiffTalk (Shen et al., 2023) was the first to modify lip movements using audio and diffusion models, but it does not extend to driving other head parts. EMO (Tian et al., 2024) was the first to leverage LDM (Rombach et al., 2022) and audio features to achieve overall motion in portraits. V-Express (Wang et al., 2024) controls the overall motion amplitude by adjusting audio attention weights, while Hallo (Xu et al., 2024) designed a hierarchical module to regulate the motion amplitude of different regions. Sonic (Ji et al., 2024) focus on global audio perception to obtain better talking head results. In summary, current audio-driven diffusion model approaches have not taken into account that each portrait should exhibit a corresponding emotion while speaking, which is essential for generating higher-quality portrait videos.
Emotional Talking Head Generation. Previous research has explored several GAN-based methods  (Ji et al., 2021; Liang et al., 2022) for extracting style information to apply in talking head generation. MEAD (Wang et al., 2020) and Emotion (Sinha et al., 2022) directly inject style labels into the network to drive the corresponding emotions. GC-AVT (Liang et al., 2022) and EAMM (Ji et al., 2022) map the facial expressions in the source video to each frame in the target video. LSF (Wu et al., 2021) and StyleTalk (Ma et al., 2023) employ 3D Morphable Models (3DMM) to extract facial information and construct style codes that drive the desired styles. EAT (Gan et al., 2023) achieves emotion control through parameter-efficient adaptations. Building on these approaches, our framework is first to disentangle identity and cooperate emotion to generate better emotion features into diffusion model and produce high-quality emotional talking portrait videos.
@@ 3. Method
Problem Formulation.  The goal of THG is to generate a talking head video under the control of a reference portrait image, audio, and emotion prior. Among these conditions, the reference portrait image provides the background and facial identity, the audio guides the overall head motion and lip movements. In addition, we propose a disentangled emotion embedder to extract the emotion prior from the visual and audio content of an emotion reference video, which is used to determine facial emotion.
@@
Overview.  As depicted in Figure 2, our DICE-Talk consists of three important designs: (1) A disentangled emotion embedder that extracts emotions from video and audio inputs, representing them as a Gaussian distribution. (2) A correlation-enhanced emotion conditioning module that stores emotion features in a bank and injects them into a diffusion model. (3) An emotion discrimination objective that improves emotion expression by classifying emotions during denoising. The diffusion process combines a reference image, audio features, emotion data, and the emotion bank to guide facial animation. Audio features come from Whisper-Tiny (Radford et al., 2023), and following (Ji et al., 2024), we use an attention layer for processing. Finally, after denoising, a VAE decoder generates expressive portrait frames with the target emotion.
@@ 3.1. Disentangled Emotion Embedder
where $\tau$ denotes a temperature parameter, $\mathcal{S}^{n}$ denotes all negative samples for $s$, and $\zeta(s_{i},s_{j})$ is the cosine similarity between sample pairs $(s_{i},s_{j})$.
@@ 3.2. Correlation-Enhanced Emotion Conditioning
Emotion Projection.  To utilize the emotion feature $E_{s}$ obtained in Sec. 3.2 and guide the denoising process, then $E_{s}$ is injected into the diffusion UNet through an additional emotion attention layer, where it interacts with other features via a cross-attention mechanism to supplement additional facial details such as expressions. The information of the emotion feature $E_{s}$ can be injected into the spatial and temporal cross-attention layers to provide spatial knowledge as follows:
@@ 4.2. Quantitative Comparison
We compare with several previous works, including SadTalker (Zhang et al., 2023), AniPortrait (Wei et al., 2024), V-Express (Wang et al., 2024), Hallo (Xu et al., 2024), Hallo2 (Cui et al., 2024), EchoMinic (Chen et al., 2024), Sonic (Ji et al., 2024), EAMM (Ji et al., 2022), StyleTalk (Ma et al., 2023), and EAT (Gan et al., 2023). Among these, EAMM, StyleTalk, and EAT are also focused on emotion modeling.
@@
As shown in Table 1, our method ranks among the top in metrics on both the MEAD and out-of-domain datasets. Our method achieved the best emotion scores, reflecting that our approach generates vivid and accurate facial expressions through emotion priors. In the quality assessment of videos, single-frame images, and lip-sync accuracy, our method is close to the current best methods, as evidenced by the FVD, FID, and SyncNet scores. The results in Table 2 further indicate that our method maintains high visual quality even on datasets with no obvious emotions like HDTF, which do not have obvious emotions. These results demonstrate that our method can ensure video generation quality while achieving diversity and accuracy in emotional expression.
@@ 4.3. Qualitative Comparison
4.4. User Study
We conducted a subjective evaluation of the results on open datasets, assessing the comparative methods along four dimensions: emotional expression, lip-sync accuracy, video quality, and video smoothness. 20 participants rated the results of five comparative methods on a scale of 1 to 5 in a total of 40 video sets. As shown in Table 3, our DICE-Talk significantly outperformed others in the dimensions of emotional expression and video smoothness, with a notable 66.5% improvement in emotional expression. Although StyleTalk and EAT achieved good results in objective metrics, they compromised the original identity and reduced visual quality while generating emotions, leading to lower subjective emotion scores. Figure 4(a) illustrates two clear examples. In terms of lip-sync accuracy and video quality, our method is very close to the state-of-the-art Sonic and far surpasses other methods that model emotions. Video results from the user study are included in the supplementary materials.